Find the value of k if the lines 2x+6y+3=0 and 3x-4y+k=0 have the same x-intercept.
Answer by ikleyn(52780)

.
1.  The line  2x + 6y + 3 = 0  has x-intercept  2x + 3 = 0,  or  x = -3/2.


2.  The line  3x - 4y + k = 0  has x-intercept  3x + k = 0,  or  x = -k/3.


3.  The two intercepts coincide if  -3/2 = -k/3,   which gives  k= 9/2 = 4.5.

Solved.